TERAHERTZ AND FAR-INFRARED SPECTROSCOPY OF HIGH-J TRANSITIONS OF THE GROUND AND v_{2} = 1 STATES OF NH_{3}

\underline{O. PIRALI}, M.-A. MARTIN, M. VERVLOET and D. BALCON, Ligne AILES–Synchrotron SOLEIL, L’Orme des Merisiers Saint-Aubin, 91192 Gif-sur-Yvette, France; S. YU, J. PEARSON and B. DROUIN, Jet Propulsion Laboratory, California Institute of Technology, Pasadena, California 91109, USA; C. P. ENDRES, I. Physikalisches Institut, Universität zu Köln, 50937 Köln, Germany; T. SHIRAISHI, K. KOBAYASHI, and F. MATSUSHIMA, Department of Physics, University of Toyama, Gofuku, Toyama 930-8555, Japan.

Since its first detection in 1968, ammonia was discovered as a major constituent of several planetary atmospheres. More recently, ammonia has been suggested in the atmosphere of cool brown dwarf ^{a} and is expected to be present in quantity in the atmospheres of many newly discovered exoplanets and brown dwarf stars where temperatures are in the order of 1000 K ^{b}. For such temperatures, spectroscopic knowledge of ammonia’s IR spectrum needs to be improved both in term of line positions and intensities. Even for the two lowest vibrational levels (ground state and v_{2} = 1) its large amplitude inversion motion complicates the spectral modelling and the experimental dataset have been (up to now) limited to low quantum numbers (J of about 20). We associated experimental results obtained from far infrared techniques^{c}^{d} and terahertz spectroscopy^{e} to obtain accurate energies for highly excited J levels (as high as J=35) in the ground state and v_{2} = 1. This work significantly increases the experimental dataset available to support astronomical observations; we will present the techniques developed in this work as well as the spectral analysis and fit of the new dataset.

aP. Delorme, X. Delfosse, L. Albert, et al., Astronomy and Astrophysics, 482, 961 (2008)

bM.J. Burgdorf, G.S. Orton, T. Encrenaz, G.R. Davis, S.D. Sidher, E. Lellouch, B.M. Swinyard, Planetary and Space Science, 52, 379 (2004)

cO. Pirali and M. Vervloet, Chemical Physics Letters, 423, 376–381 (2006)

dF. Matsushima, H. Odashima, D. Wang, S. Tsunekawa, and K. Takagi, Jpn. J. Appl. Phys. 33, 315 (1994)

eB.J. Drouin, F.W. Maiwald, and J.C. Pearson, Rev. Sci. Instrum., 76, 093113 (2005)